OpenStudy (anonymous):

which of the following is an identity? A) sin^2x sec^2x+1=cot^2x csc^2x B) sin^2x-cos^2x=1 C) (cscx+cot)^2=1 D) csc^2x+cot^2x=1

OpenStudy (anonymous):

i think its b)

OpenStudy (lgbasallote):

it's not b) the identity should be \(\sin^2 x + \cos^2 x = 1\)

OpenStudy (anonymous):

b is not correct. sin^2x + cos^2x = 1

OpenStudy (anonymous):

I wrote B correctly..that is why I'm confused

OpenStudy (lgbasallote):

another identity you should know \(1 + \cot^2 x = \csc^2 x\)

OpenStudy (lgbasallote):

therefore \(\csc^2 x - \cot^2 x = 1\)

OpenStudy (lgbasallote):

are you following my logic @mInnocenti ? do you get what im implying?

OpenStudy (anonymous):

I'm following. thank you :)

OpenStudy (lgbasallote):

so do you know the correct answer now?

OpenStudy (anonymous):

b)

OpenStudy (anonymous):

Hold on, option C is not the same as \[\csc^2x−\cot^2x=1\] \[C= \csc^2x + 2 cotxcscx + \cot^2x\]

OpenStudy (lgbasallote):

who said it was c? lol

OpenStudy (anonymous):

Oh I hadn't seen that option....lol never mind.

OpenStudy (anonymous):

I thik it's b? lol

OpenStudy (anonymous):

b is not correct as stated earlier.

OpenStudy (anonymous):

Look at what @lgbasallote told you. B is incorrect.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

lgbasallote also told you that csc^2x - cot^2x =1 which implies csc^2x + cot^2x cannot equal 1 as well...

OpenStudy (anonymous):

right...I think I'm too weary to be of help right now, lol But remember also \[\sec^2x-1=\tan^2x\]

OpenStudy (anonymous):

only A is left.. lol

OpenStudy (anonymous):

Thank you everyone!

OpenStudy (lgbasallote):

exactly lol